During 2020, Monty Furniture Limited purchased a railway carload of wicker chairs. The manufacturer of the chairs sold them to Monty for a lump sum of $58,600, because it was discontinuing manufacturing operations and wanted to dispose of its entire stock. Three types of chairs are included in the carload. The three types and the estimated selling price for each are as follows:

Type No. of Chairs Estimated Selling
Price per Chair
Lounge chairs 470 $93
Straight chairs 640 53
Armchairs 370 83


Monty estimates that the costs to sell this inventory would be $4 per chair. During 2020, Monty sells 390 lounge chairs, 270 armchairs, and 150 straight chairs, all at the same prices as estimated. At December 31, 2020, the remaining chairs were put on sale: the lounge chairs at 25% off the regular price, the armchairs at 30% off, and the straight chairs at 40% off. All were expected to be sold at these prices.

A) What is the total cost of the chairs remaining in inventory at the end of 2020, using the relative sales value method?

B) What is the net realizable value of the chairs remaining in inventory?

C) What is the appropriate inventory value to be reported on the December 31, 2020 statement of financial position, assuming the lower of cost and NRV is applied on an individual item basis?
Answer #1

A) What is the total cost of the chairs remaining in inventory at the end of 2020, using the relative sales value method?

Answer:

Total Costs 22,545.40

Calculation:

(Amounts and percentages are rounded)

The formulas are given along with the table headings.

Type Number of Units (a) Selling Price (b)

Total

(c) = (a) * (b)

%

(d)

Allocation (e)= 58,600 * (d) Each (f) =(e)/(a) Ending Inventory (g) Cost of Inventory (f) * (g)
Lounge 470 93 43,710 40.3% 23,616 50.25 80 4,020.00
Straight 640 53 33,920 31.3% 18,342 28.66 490 14,043.40
Arm 370 83 30,710 28.3% 16,584 44.82 100 4,482.00
108,340 22,545.40

% calculation:

Type Total/Overall Total =

Lounge = 43,710/108,340 = 40.3%

Straight = 33,920/108,340 = 31.3%

Arm = 30,710/108,340 = 28.3%

Ending Inventory :

Lounge = 470 - 390 = 80

Straight = 640 - 150 = 490

Arm = 370 - 270 = 100

B) What is the net realizable value of the chairs remaining in inventory?

Answer:

Net Realizable Value 24,292

Calculation:
(Amounts and percentages are rounded)

Type Selling Price (a) Discount (b) New Price (c)= (a)-(a)*(b) Less Selling Cost (d) Each
(e)=(c)-(d)
Ending Inventory (f) Net Realizable Inventory
(e) *(f)
Lounge 93 25% 69.750 4 65.75 80 5,260
Straight 53 40% 31.800 4 27.8 490 13,622
Arm 83 30% 58.100 4 54.1 100 5,410

5,260 +13 622 + 5,410 = 24,292

C) What is the appropriate inventory value to be reported on the December 31, 2020 statement of financial position, assuming the lower of cost and NRV is applied on an individual item basis?

Answer:

Inventory Value 22,124

Calculation:

(Amounts and percentages are rounded)


Lower of cost and NRV is applied. So,

We need to take Lower of cost or NRV which is applicable to each.

Lounge = 4,020 (Cost)

Straight = 13,622 (NRV)

Arm = 4,482 (Cost)

And we need to total these amounts to get the appropriate inventory value to be reported on the December 31, 2020 statement of financial position

Hence, 4,020 + 13,622 + 4,482 = 22,124